ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 143, 2004

Buhovsky, L. Homology of Lagrangian submanifolds in cotangent bundles

181

Gil-Medrano, O., J. C. González-Dávila and L. Vanhecke Harmonicity and minimality of oriented distributions

253

González-Dávila, J. C. See O. Gil-Medrano

Hájek, P. and M. Johanis Isomorphic embeddings and harmonic behaviour of smooth operators

299

Humphries, S. P. Finite Hurwitz braid group actions for Artin groups 189

Johanis, M. See P. Hájek

Kang, D. and S. Peng Existence of solutions for elliptic problems with critical Sobolev–Hardy exponents

281

Kellerhals, R. On the structure of hyperbolic manifolds  361

Leung, D. H. and W.-K. Tang ℓ^{1}-spreading models in mixed Tsirelson space

223

Li, B. Q. Entire solutions of certain partial differential equations in

C^{n} 131

Lichtman, A. I. and D. S. Passman Finitely generated simple algebras: A question of B. I. Plotkin

341

Loehr, N. A., G. S. Warrington and H. S. Wilf The combinatorics of a three-line circulant determinant

141

Lusky, W. On Banach spaces with unconditional bases  239

Passman, D. S. See A. I. Lichtman

Peng, S. See D. Kang

Sela, Z. Diophantine geometry over groups IV: An iterative procedure for validation of a sentence

1

Stanley, R. P. Recent developments in algebraic combinatorics  317
Talagrand, M. Type and infratype in symmetric sequence spaces  157

Tang, W.-K. See D. H. Leung

Vanhecke, L. See O. Gil-Medrano

Warrington, G. S. See N. A. Loehr

Wilf, H. S. See N. A. Loehr